OpenStudy (anonymous):

please help The 11th term of an arithmetic progression is 1. the sum of the first 10 terms is 120. find the 4th term :S formula for the... sum of terms=1/2n(a+l)=1/2n{2a+(n-1)d} thanks

OpenStudy (yuki):

okay

OpenStudy (yuki):

\[{11(a_1 + a_{11}) \over 2} = 120 + 1 \]

OpenStudy (yuki):

so\[{a_1 + 1} = 22\] thus a_1 = 21

OpenStudy (yuki):

Then you can say\[a_{11} = a_1 + 10d\] \[a_11 = 1 = 21 + 10d\]

OpenStudy (yuki):

so d = -2

OpenStudy (yuki):

that means that the 4th term is 21+3(-2) = 15

OpenStudy (yuki):

:)

OpenStudy (yuki):

did it make sense ?

OpenStudy (anonymous):

but you didnt even use 4 once :S??? and i dont really understand where 21+3(-2)=15 comes from :(

OpenStudy (yuki):

ok, \[a_n = a_1 + d(n-1)\] so\[a_4 = a_1 + (-2)(4-1) = 21 + 3(-2)\]

OpenStudy (yuki):

the idea of arithmetic sequence is that you are "Adding the same number for all sequence"

OpenStudy (anonymous):

thanks :)

OpenStudy (yuki):

np :)

OpenStudy (anonymous):

sorry but i first thought that i knew where 11(a1+a11)2=120+1 came from but i dont :S

OpenStudy (yuki):

Hope you read this.

OpenStudy (yuki):

the first ten numbers add up to 120, so 10(a1+a10) /2= 120 now if you add a11 to it, it is 120 + 1 but adding a11 to the sum from a1 to a10 is as same as adding from a1 to a11 so 11(a1 + a11)/2 = 121